# Meta-Transfer Derm-Diagnosis: Exploring Few-Shot Learning and Transfer Learning for Skin Disease Classification in Long-Tail Distribution

### v1

@@ Abstract
Addressing the challenges of rare diseases is difficult, especially with the limited number of reference images and a small patient population. This is more evident in rare skin diseases, where we encounter long-tailed data distributions that make it difficult to develop unbiased and broadly effective models. The diverse ways in which image datasets are gathered and their distinct purposes also add to these challenges. Our study conducts a detailed examination of the benefits and drawbacks of episodic and conventional training methodologies, adopting a few-shot learning approach alongside transfer learning. We evaluated our models using the ISIC2018, Derm7pt, and SD-198 datasets. With minimal labeled examples, our models showed substantial information gains and better performance compared to previously trained models. Our research emphasizes the improved ability to represent features in DenseNet121 and MobileNetV2 models, achieved by using pre-trained models on ImageNet to increase similarities within classes. Moreover, our experiments, ranging from 2-way to 5-way classifications with up to 10 examples, showed a growing success rate for traditional transfer learning methods as the number of examples increased. The addition of data augmentation techniques significantly improved our transfer learning based model performance, leading to higher performances than existing methods, especially in the SD-198 and ISIC2018 datasets. All source code related to this work will be made publicly available soon at the provided URL.
Keywords Few-shot learning $\cdot$ Long-tail distribution  $\cdot$ Medical image classification  $\cdot$ Skin disease classification  $\cdot$ Transfer learning
@@ 1 Introduction
Over the past decade, the field of medical image analysis has witnessed remarkable advancements, primarily driven by the development of deep convolutional neural networks and the availability of extensive labeled image datasets. These advancements have notably impacted various tasks, including organ segmentation [1, 2], tumor segmentation [3, 4], and disease detection [5, 6]. Although abundant data exists for common diseases, a significant gap persists in data availability for the over 6,000 known rare diseases, affecting approximately 7% of the global population [7]. The diagnosis of these rare diseases, including some skin conditions, presents unique challenges, particularly due to the limited number of clinical examples available for training deep learning models. The automatic classification of skin lesions exemplifies these challenges, as it is complicated by the long-tailed distribution of skin disease datasets, the subtle variations in lesion appearances, and the overall scarcity of sufficient image data [8].
@@
Figure 1: The figure illustrates the distributions of the datasets used in this study, namely SD-198[9], Derm7pt[10] and ISIC2018[11] . The observed data distribution exhibits a long-tailed nature, and an examination reveals that some classes have very few instances. The base classes (common diseases) used during the training process include the train (green) and validation (yellow) classes. Novel classes (new/rare diseases) are indicated with the test label (red). In the Derm7Pt dataset, classes with a significantly low number of examples are displayed as deleted (grey) and deemed unusable.
@@
Various studies have been conducted to address the problem of skin disease classification using deep learning approaches. Recent advancements in this field are mainly in three categories: methods based on transfer learning [12, 13], those relying on few-shot learning [14, 15, 8, 16, 17, 18, 19], and approaches using cross-domain few-shot learning [20]. The state of the art models in this domain, such as Meta-DermDiagnosis, MetaMed, and PCN models [8, 18, 14], are designed to extract and learn high-level, domain-specific features during their training process.
@@
In their study on the ISIC2018 dataset, Li et al. (2020) introduced the Difficulty Aware Meta-Learning (DAML) model, recognizing that randomly selected tasks vary in difficulty [15]. They proposed optimizing the meta-optimization process by dynamically assessing the significance of challenging tasks. Similarly, Mahajan et al. (2020) developed the Meta-DermDiagnosis model, experimenting with the SD-198, Derm7pt, and ISIC2018 datasets [8]. This model innovatively replaced traditional convolutional layers in Prototypical Networks and Reptile with Group Equivariant Convolutions, enhancing resistance to data transformations like symmetry and rotation, thus yielding robust features. However, this method received critique for its reliance on datasets with symmetric orientation. On a related note, Singh et al. (2021) highlighted the efficacy of MixUp, CutOut, and CutMix as data augmentation techniques in the medical field when integrated with the MAML algorithm in their MetaMed model, particularly on the ISIC2018 dataset [18]. Both MetaMed and Meta-DermDiagnosis models aim to enhance feature representation by broadening data augmentation and diversification. Further, Prabhu et al. (2019) proposed the Prototypical Clustering Networks (PCN) model, designed to effectively handle intra-class variability in few-shot learning scenarios [14].
@@
The studies previously discussed have employed episodic learning as a means to acquire knowledge that can be transferred to new classes. This approach involves the arrangement of learning problems into episodes, each composed of small training and validation subsets, designed to simulate the scenarios encountered during evaluation. However, [21] criticized this method, arguing that the constraints of episodic learning are unnecessary and that using training groups in this manner is data-inefficient. Similarly, [22] argued that in tasks involving rare classes, the effectiveness of rapid adaptation is more attributable to the quality of the learned representation than to the few-shot learning algorithm itself. They incorporated self-supervised learning in their model to enhance this representation. In a similar context, [23] proposed the concept of Meta-transfer learning, which is aimed at refining the learned representation. They conducted various experiments to compare the effectiveness of fine-tuning pre-trained Deep Neural Network (DNN) models with and without episodic learning. Their findings underscore the significance of integrating transfer learning with few-shot learning methodologies.
@@
The existing literature consistently highlights the ongoing challenge of effectively managing the issue of long-tailed data distribution in skin disease studies [8, 24, 25]. Although current methods show promise in specific dataset contexts, their ability to generalize broadly remains limited. Most research in this field has been directed towards developing dataset-specific solutions, addressing the unique challenges and sensitivities arising from variations in class counts, data formats, and other critical characteristics.
In our study, we aim to conduct an in-depth analysis of the benefits and drawbacks of integrating few-shot learning, episodic learning, and transfer learning. Our main focus is on developing a foundational framework specifically designed to utilize the inherent properties of long-tailed skin datasets. While acknowledging the potential benefits of performance-enhancing methods, our base approach distinctly diverges from such techniques.
In this context, we implemented four distinct training methods, each evaluated using consistent metrics for newly introduced classes. Our first method, Few-Shot Episodic Transfer Learning (FETL), involves adapting the model for the dataset through fine-tuning weights that were initially pre-trained on ImageNet, coupled with episodic few-shot learning. The second strategy, Few-Shot Episodic Learning (FEL), diverges from the first by employing episodic learning on models without any pre-training. The third approach, Deep Transfer Learning (DTL), focuses on fine-tuning models using deep neural networks (DNN) with pre-trained ImageNet weights, but notably omits episodic learning. Lastly, our fourth method, Deep Learning (DL), acts as a baseline, utilizing pre-trained ImageNet weights without additional adjustments or fine-tuning. Following the establishment of our approach, we aimed to enhance performance by incorporating well-known data augmentation techniques such as CutMix, MixUp, and ResizeMix into the DTL model.
@@
To evaluate the effectiveness of our proposed methodologies, we carried out extensive testing across three benchmark skin disease datasets: SD-198, Derm7pt, and ISIC2018. This comprehensive analysis provided us with critical insights into the most effective strategies for tackling the challenge of long-tailed data distribution in skin diseases. The key contributions of our study are summarized as follows:
@@
  • We introduce a novel methodology designed to evaluate various model training approaches, using a consistent benchmark test set specifically made for long-tail distributions in rare skin diseases. This evaluation is conducted through episodic testing. To the best of our knowledge, this is the first time such a thorough methodological analysis has been conducted in this domain.
@@
  • We demonstrated that combining transfer learning with few-shot learning significantly enhances both the learned representation and the testing performance in the context of rare skin diseases. Using our proposed model based on transfer learning, along with augmentation techniques like MixUp, CutMix, and ResizeMix, has helped us surpass state-of-the-art results in some settings on the SD198 and ISIC2018 datasets.
@@ 2.1 Transfer Learning for Medical Image Domain
In deep learning, a powerful transfer learning method involves adapting a pre-trained model for a new task, commonly referred to as fine-tuning (FT). Models that have been pre-trained on extensive datasets have demonstrated superior generalization performance compared to models initialized randomly [26]. Various techniques are employed to facilitate the transfer of knowledge between different source-target domains [27, 28, 29, 30].
Transfer learning (TL) is frequently applied in addressing the classification problem of skin diseases. For instance, [31] aimed to improve performance on the ISIC2018 dataset by using ISIC2016 and ISIC2017 as source datasets and EfficientNetB0/B1, SeReNext50 as backboned models. [32] performed transfer from the ImageNet dataset to the ISIC2018 dataset using ResNet50. Similarly, [33] employed ImageNet, ResNet, and DenseNet architectures for transfer to the ISIC2017 dataset. [34], with a similar objective, experimented with the VGG16 architecture. A common inference drawn from these studies is that the use of transfer learning, particularly when ImageNet or various dermatological datasets are used as source datasets and supported by architectures like ResNet or DenseNet, contributes significantly to addressing this problem. A more comprehensive literature discussion is provided in [35].
In addition to traditional learning methods, transfer learning has also been explored in the Few-Shot Learning (FSL) domain for the classification problem of skin diseases. For instance, the MetaMed study [18], working with FSL algorithms on the ISIC2018 dataset, compared and interpreted the results of models trained through TL with their proposed method. In this process, they employed a shallow-layered architecture in the TL model and conducted training using only base classes. In contrast, as indicated in the literature, we adapted the MobileNet and DenseNet architectures to the Few-Shot Learning domain by utilizing pre-trained weights from ImageNet. We discussed detailed comparison results in Section 5.3.
@@ 2.3 Few-Shot Learning for Skin Disease Classification
In the context of transfer-learning-based algorithms, [25] proposed a model named PFEMed, suggesting a dual-encoder structure. This brings about one encoder with fixed weights pre-trained on large-scale public image classification datasets and another encoder trained on the target medical dataset. On the other hand, [24] designed a dual-branch framework and improved performance using a model employing prototypical networks and contrastive loss. To address the challenge of observing diverse subgroups within dermatological disease clusters, [19] introduced the Sub-Cluster-Aware Network (SCAN) model. SCAN utilizes a dual-branch structure to enhance feature explanation, learning both class-specific features for disease differentiation and subgroup-related features.
In this study, we integrate Prototypical Networks from Few-Shot Learning (FSL) algorithms with transfer learning techniques. This fusion involves adapting MobileNetV2 and DenseNet121 backbone architectures, coupled with pretrained ImageNet weights, specifically to tackle the challenge of rare skin disease identification. Our experimental analysis highlights the limitations of episodic training in optimally leveraging data. Central to our approach is the establishment of a foundational framework that relies solely on the inherent properties of the long tailed datasets, without the necessity for supplementary data. This includes strategies aimed at enhancing model performance, such as extensive data augmentation techniques like MixUp, CutMix, and ResizeMix. The utilization of these techniques contributes to achieving our research objectives.
@@ 3 Datasets and Evaluation
The SD-198 dataset [9] comprises 198 detailed categories of skin diseases, including eczema, acne, rosacea, and various cancer conditions. These categories contain 6584 clinical images contributed by patients and dermatologists, showcasing diverse characteristics like color, exposure, lighting, and size. The dataset captures a wide range of patients in terms of age, gender, disease location, skin color, and disease stage. Originally divided into a 50% training set and a 50% test set, the images are captured at 1640 × 1130 pixels using digital cameras or mobile phones. To align with the comparison criteria set by Meta-DermDiagnosis [8] and SCAN [19], we resized all data to 224 x 224 pixels. For testing, we focused on 70 classes representing rare diseases, each having fewer than 20 images, while the remaining 128 classes were used for training. The dataset distribution is visualized in Figure 1, and sample images are provided in Figure 2-a.
The Derm7pt dataset [10] comprises over 2,000 clinical and dermoscopy images grouped into 20 distinct classes. This dataset, comprising both clinical and dermoscopic images, provides predictions based on a 7-point checklist for the malignancy of skin lesions, making it suitable for training and evaluating computer-aided diagnosis (CAD) systems. The original images have dimensions of 768 × 512 pixels; however, for the purpose of our experimental studies, they have been resized to 224 × 224 pixels. To facilitate the comparison of our experimental results with the Meta-DermDiagnosis and SCAN study, we adopted similar train-test set differentiations. Within the Derm7pt dataset, two categories are excluded from our experiments: ’miscellaneous’ (encompassing unspecified skin diseases) and ’melanoma’ (due to its solitary instance, preventing a train-test division). Among the 18 lesion categories in this dataset, 13 classes are allocated for training, while the remaining categories are reserved for testing. The novel set consists of 5 classes, with each class containing 10 to 34 images. This distinction strategy aims to mimic the ability to generalize to infrequent skin diseases by placing classes with limited data in the test set. For visual reference, selected examples of skin lesion images can be found in Figure 2-b. The dataset distribution is also shared in Figure 1.
The ISIC 2018 Skin Lesion dataset [11] comprises 10,015 dermoscopic images that are categorized by expert pathologists into seven distinct skin lesion classes. Within this dataset, 7,515 images are allocated to the training set, while the remaining 2,500 images constitute the test set, following a standardized partition. Dermoscopic images often position the target lesion at the center. We adhered to similar scaling and data partitions as the Meta-DermDiagnosis[8] and PFEMed [25] study for our experiments. Consequently, the image resizing process transforms images from 600 × 450 pixels to 224 x 224 pixels, and a subset comprising four base classes and three novel classes is selected to form few-shot classification tasks. Refer to Figure 2-c for exemplar images drawn from the dataset. For data distribution, please see Figure 1.
@@
Figure 3: Overall framework of our pipeline: Meta-Transfer Derm-Diagnosis. a) Episodic learning is combined with DenseNet and MobileNet architectures without the use of ImageNet weights. b) ImageNet pre-trained weights are utilized along with the application of an episodic learning strategy. c) Pre-trained weights and all base class data are employed for fine-tuning with ImageNet. d) Only ImageNet weights are utilized without fine-tuning. e) Detailed diagram illustrating the use of episodic learning and prototypical networks. It is applied in the continuation of parts a and b. f) Common evaluation scheme using novel data across segments a, b, c, and d.
@@
Figure 4: The flowchart illustrating the components and the training strategies of the FEL, FETL, DTL, and DL models within the proposed framework.

### v2

# Zeynep Özdemir*

@@ Abstract
Building accurate models for rare skin diseases remains challenging due to the lack of sufficient labeled data and the inherently long-tailed distribution of available samples. These issues are further complicated by inconsistencies in how datasets are collected and their varying objectives. To address these challenges, we compare three learning strategies: episodic learning, supervised transfer learning, and contrastive self-supervised pretraining, within a few-shot learning framework. We evaluate five training setups on three benchmark datasets: ISIC2018, Derm7pt, and SD-198. Our findings show that traditional transfer learning approaches, particularly those based on MobileNetV2 and Vision Transformer (ViT) architectures, consistently outperform episodic and self-supervised methods as the number of training examples increases. When combined with batch-level data augmentation techniques such as MixUp, CutMix, and ResizeMix, these models achieve state-of-the-art performance on the SD-198 and Derm7pt datasets, and deliver highly competitive results on ISIC2018. All the source codes related to this work will be made publicly available soon at the provided URL.
Keywords Few-shot learning $\cdot$ Long-tail distribution $\cdot$ Medical image classification $\cdot$ Skin disease classification $\cdot$ Self-supervised learning $\cdot$ Transfer learning $\cdot$ Explainability $\cdot$ Uncertainty estimation
@@ 1 Introduction
Figure 1: Skin lesion classification framework: (a) benchmark datasets, (b) transfer learning strategies.
@@
Over the past decade, the field of medical image analysis has witnessed remarkable advancements, primarily driven by the development of deep convolutional neural networks and the availability of extensive labeled image datasets. These advancements have notably impacted various tasks, including organ segmentation [1, 2], tumor segmentation [3], and disease detection [4]. Although abundant data exists for common diseases, a significant gap persists in data availability for the over 6,000 known rare diseases, affecting approximately 7% of the global population [5]. The diagnosis of these rare diseases, particularly certain skin conditions, presents unique challenges specific to the domain. Skin disease datasets, unlike those for natural image classification, often exhibit long-tailed distributions, where a few common classes dominate while rare classes are underrepresented. Additionally, variations in image quality and the presence of artifacts such as hair, rulers, and ink markings further complicate the classification process. The diversity of imaging modalities, including clinical and dermoscopic images, adds another layer of complexity, as these modalities vary in resolution, feature representation, and clinical relevance. Anatomical and biological diversity, subtle visual distinctions, and patient demographics introduce further complexities that directly impact model performance. Moreover, the need for expert clinicians to annotate datasets and strict data privacy regulations that limit data sharing represent significant obstacles in this field. These challenges underscore the necessity for developing advanced, domain-specific methodologies designed to address the complexities of skin disease classification [6, 7, 8, 9].
@@
Figure 2: The figure shows the distributions of the datasets SD-198[10], Derm7pt[11], and ISIC2018[12], highlighting their long-tailed nature with some classes having very few instances. Base classes (common diseases) are marked as train (green) and validation (yellow), while novel classes (rare diseases) are labeled as test (red). In the Derm7pt dataset, classes with very few examples are shown as deleted (grey) and excluded from use.
Given these limitations, Few-shot learning (FSL) has emerged as a promising approach to address the challenges of class imbalance and data scarcity in medical imaging. FSL offers versatile solutions in image classification, detection, and segmentation tasks across both natural and medical images [13, 14, 15, 16, 17, 18, 19, 20, 21]. Specifically, FSL methods have demonstrated notable effectiveness in navigating the complexities of skin disease classification [9].
@@
Building upon these advancements, various studies have been conducted to address the problem of skin disease classification using deep learning approaches. Recent advancements in this field are mainly in three categories: methods based on transfer learning [22, 23], those relying on few-shot learning [24, 25, 9, 26, 27, 28, 29], and approaches using cross-domain few-shot learning [30]. The state of the art models in this domain, such as Meta-DermDiagnosis, MetaMed, and PCN models [9, 28, 24] are designed to extract and learn high-level, domain-specific features during their training process.
@@
In their study on the ISIC2018 dataset, Li et al. (2020) introduced the Difficulty Aware Meta-Learning (DAML) model, addressing task variability challenges [25]. Similarly, Mahajan et al. (2020) developed the Meta-DermDiagnosis model, experimenting with the SD-198, Derm7pt, and ISIC2018 datasets [9]. This model incorporated Prototypical Networks and Reptile to enhance robustness but faced criticism for its reliance on symmetric dataset orientations. Singh et al. (2021) highlighted the efficacy of MixUp, CutOut, and CutMix as data augmentation techniques in the medical field when integrated with the MAML algorithm in their MetaMed model, particularly on the ISIC2018 dataset [28]. Both MetaMed and Meta-DermDiagnosis models aim to enhance feature representation by broadening data augmentation and diversification.
@@
The mentioned studies in the field of skin disease classification have employed episodic learning as a method to acquire knowledge transferable to new classes. This approach involves dividing learning problems into small training and validation subsets to simulate scenarios encountered during evaluation. However, some studies in the FSL domain working with natural images have criticized episodic learning, arguing that its constraints are unnecessary and that using training groups in this manner is data-inefficient [31]. Similarly, [32] argued that, in tasks involving rare classes, the effectiveness of rapid adaptation depends more on the quality of the learned representation than on the few-shot learning algorithm itself. To enhance representation quality, they incorporated self-supervised learning components into their models. In a similar context, [33] proposed the concept of Meta-Transfer Learning, which aims to refine the learned representation. These studies conducted various experiments comparing the effectiveness of fine-tuning pre-trained Deep Neural Network (DNN) models with and without episodic learning. Their findings highlight the significance of integrating transfer learning with few-shot learning methodologies.
@@
The existing literature consistently highlights the ongoing challenge of effectively managing the issue of long-tailed data distribution in skin disease studies [9, 34, 35]. Although current methods show promise in specific dataset contexts, their ability to generalize broadly remains limited. Most research in this field has focused on developing dataset-specific solutions, addressing the unique challenges and sensitivities arising from variations in class counts, data formats, and other critical characteristics. On the other hand, episodic training techniques, which have been criticized for their inefficiency in the natural image domain of FSL, continue to be utilized in this context.
In this study, we aim to explore the effects of episodic learning, transfer learning and self-supervised learning for rare skin disease classification. One of our objective was to evaluate different model training approaches, using a benchmark test set specifically for rare skin diseases. In addition to transfer learning approaches using supervised ImageNet-pretrained models such as ResNet50 [36], DenseNet121 [37], MobileNetV2 [38], Vision Transformer (ViT) [39], ConvNeXt [40], and EfficientNet [41], we extended our evaluation to include models pretrained via contrastive self-supervised learning, including SimCLR [42], MoCo-v3 [43], DINO-v2 [44], SimSiam [45], SwAV [46], and ViT-MMF [47]. Additionally, we investigated the potential of transformer-based foundation models developed for medical imaging, such as BioViL [48] and MedCLIP [49].
In this context, we implemented five distinct training approaches, each evaluated using consistent metrics for novel classes (Figure 1). Our first approach, Few-Shot Episodic Learning (FEL), applies episodic training to models initialized with random weights (i.e., without any pre-training). The second, Few-Shot Episodic Transfer Learning (FETL), extends this by incorporating ImageNet-pretrained weights, which are fine-tuned on the long-tail skin disease dataset within an episodic learning setup. The third method, Deep Transfer Learning (DTL), omits episodic training and instead fine-tunes ImageNet-pretrained models directly on base classes. The fourth, Deep Transfer Learning with Self-Supervised Learning (DTLS), uses contrastively pretrained models, which are then fine-tuned with supervised training on the target dataset. Finally, Deep Learning (DL) serves as a baseline, where ImageNet-pretrained models are evaluated without any additional training or fine-tuning.
FETL, DTL, and DTLS were designed to incorprate FSL natural image domain advancements to address rare challenge. Additionally, performance improvements were targeted by integrating data augmentation techniques, such as MixUp, CutMix, and ResizeMix, specifically adapted for the unique properties of skin disease datasets.
@@
To evaluate the effectiveness of explored methods we carried out extensive testing across three benchmark skin disease datasets: SD-198, Derm7pt, and ISIC2018. Our comprehensive analysis provide insights for effective strategies in tackling the challenge of long-tailed data distribution in skin diseases. The key contributions of our study are summarized as follows:
@@
  • We evaluate different model training approaches, using a benchmark test set specifically curated for long-tail distributions in rare skin diseases. The final comparison is conducted through episodic testing. To the best of our knowledge, this is the first time such a thorough methodological analysis has been conducted in this domain.rare skin disease.
@@
  • We systematically evaluated a wide range of pretrained and foundation models (e.g., ConvNeXt, ViT, ResNet50, MedCLIP, BioVIL) under few-shot settings, both with and without additional fine-tuning. Our results provide valuable insights into selecting strong initialization points for rare skin disease classification, emphasizing the importance of pretraining choices in low-data regimes.
  • We demonstrated that combining transfer learning and self-supervised learning with few-shot learning significantly enhances both the learned representation and the testing performance in the context of rare skin diseases.
@@ 2.1 Transfer Learning for Medical Image Domain
In deep learning, transfer learning involves adapting a pre-trained model for a new task, commonly referred to as fine-tuning (FT). Models pre-trained on large datasets have shown superior generalization compared to randomly initialized ones [50]. Various techniques facilitate the transfer of knowledge between source and target domains [51, 52, 53, 54].
For skin disease classification, transfer learning is widely used. Architectures such as EfficientNet, ResNet, and DenseNet, typically pre-trained on ImageNet or dermatological datasets, have been transferred to datasets like ISIC2017 and ISIC2018 [55, 56, 57, 58], demonstrating notable performance gains in data-scarce scenarios. A more comprehensive discussion is provided in [59].
Beyond conventional supervised pretraining, recent advances in self-supervised learning (SSL) have introduced novel mechanisms for learning transferable and robust representations without labels. In few-shot skin disease classification, studies such as MetaMed [28] explored episodic training with data augmentation, using shallow architectures and transfer learning on ISIC2018. In contrast, our approach leverages deeper architectures to enhance feature representation, aiming to better capture the complexity of rare skin disease patterns.
@@ 2.3 Few-Shot Learning for Skin Disease Classification
In the context of transfer-learning-based algorithms, the study in[35] proposed a model named PFEMed, suggesting a dual-encoder structure. This brings about one encoder with fixed weights pre-trained on large-scale public image classification datasets and another encoder trained on the target medical dataset. On the other hand, [34] designed a dual-branch framework and improved performance using a model employing prototypical networks and contrastive loss. To address the challenge of observing diverse subgroups within dermatological disease clusters, [29] introduced the Sub-Cluster-Aware Network (SCAN) model. SCAN utilizes a dual-branch structure to enhance feature explanation, learning both class-specific features for disease differentiation and subgroup-related features.
Building upon previous studies, we present a comprehensive framework for few-shot rare skin disease classification by systematically evaluating a wide range of backbone architectures and training strategies. Specifically, we experimented with architectures including ResNet50 [36], DenseNet121 [37], MobileNetV2 [38], ViT [39], ConvNeXt [40], and EfficientNet [41], most of which were pretrained using supervised learning on ImageNet. To explore alternative pretraining paradigms, we also employed contrastive self-supervised methods such as SimCLR [42], MoCo-v3 [43], SwAV [46], and SimSiam [45], as well as distillation-based approaches like DINOv2 [44] and ViT-MMF [47], particularly on ResNet50 and ViT backbones. Additionally, transformer-based foundation models tailored for medical imaging, including BioViL [48] and MedCLIP [49], were included as part of our extended transfer learning evaluation.
Our comparative analysis focuses on three major training paradigms: traditional supervised transfer learning, contrastive self-supervised pretraining, and episodic few-shot learning. While contrastive pretraining showed benefits in certain cases, our results indicate that supervised transfer learning generally yields more robust performance in few-shot settings. Furthermore, episodic learning methods often failed to fully exploit the available data. To improve generalization, we incorporated augmentation techniques such as MixUp, CutMix, and ResizeMix, customized to address the characteristics of dermatological datasets. Importantly, our evaluations rely solely on existing benchmark datasets and do not require any supplementary external data, making the findings potentially applicable to other low-resource rare disease classification tasks. Detailed results and comparative insights are presented in Section 5.4.
@@ 3 Datasets and Evaluation
All datasets used in this study were standardized to an input size of 224 $\times$ 224 pixels for consistency with prior works [9, 29, 35]. Their distributions are presented in Figure 2, while representative examples are shown in Figure 3.
The SD-198 dataset [10] comprises 6,584 clinical images across 198 categories (e.g., eczema, acne, rosacea, and skin cancers). Captured using digital cameras or mobile phones at 1640 $\times$ 1130 resolution, the images exhibit broad diversity in patient age, gender, skin type, disease stage, and imaging conditions. Following previous works, 70 rare classes (fewer than 20 samples) were designated for testing, while the remaining 128 classes served for training.
The Derm7pt dataset [11] contains over 2,000 clinical and dermoscopic images grouped into 20 classes, annotated using a 7-point checklist for malignancy. Two categories miscellaneous (unspecified diseases) and melanoma (single case) were excluded, leaving 18 categories, of which 13 were used for training and 5 for testing (10–34 samples per novel class).
The ISIC 2018 dataset [12] includes 10,015 dermoscopic images across seven lesion classes labeled by expert pathologists. Following the standard split, 7,515 images form the training set and 2,500 form the test set. In line with prior studies, a subset comprising four base classes and three novel classes was selected for few-shot classification tasks.
@@
Figure 4: Overall framework of our pipeline: Meta-Transfer Derm-Diagnosis. a) Episodic learning is combined with DenseNet and MobileNet architectures without the use of ImageNet weights. b) ImageNet pre-trained weights are utilized along with the application of an episodic learning strategy. c) Pre-trained weights and all base class data are employed for fine-tuning with ImageNet. d) Pre-trained weights and all base class data are employed for self-supervised training with ImageNet. e) Only ImageNet weights are utilized without fine-tuning. e) Detailed diagram illustrating the use of episodic learning and prototypical networks. It is applied in the continuation of parts a and b. g) Common evaluation scheme using novel data across segments a, b, c, d and e.
@@
Figure 5: The flowchart illustrating the components and the training strategies of the FEL, FETL, DTL, DTLS and DL models within the proposed framework.